The length of a rectangle is 1 cm more than 4 times its width. If the area of the rectangle is 65 cm2, find the dimensions of the rectangle?
Answer by stanbon(75887)

The length of a rectangle is 1 cm more than 4 times its width. If the area of the rectangle is 65 cm2, find the dimensions of the rectangle?
-------------
Let the width be "w"
Then the length is "4w+1"
-------------------------
Area = width*length
65 = w(4w+1)
65 = 4w^2+w
----------
Rearrange:
4w^2+w-65 = 0
w = [-1 +- sqrt(1 - 4*4*-65)]/8
w = [-1 +- sqrt(1041)]/
w = [-1 +- 32.2645...]/8
w = 3.908...cm (width of the rectangle)
4w+1 = 16.6322...cm (length of the rectangle)
------------
Cheers,
Stan H.